Calculate how thick layer of nickel is formed on a metal plate when a current of 100 m A is lead 30 mins through a solution containing Ni2+-ions. The area of the plate is 0.10 m2.
1
Expert's answer
2012-04-19T07:39:38-0400
m = Mw*I*t/n*F

m - mass
Mw - molecular weight
n - the numberof electronsinvolved in the process
I - electric current
F - Faraday constant
t - time

m = 58.69*0.1*(30*60)/2*96485.3 = 0.05475

density is 8.908 g/cm3

V = 0.05475/8.908 = 0.0061456 cm3
h = 0.0061456/0.1 = 0.061456 cm or 0.61456 mm